OpenStudy (anonymous):

If y = (u-3)/ (u+3) and u = 8/ sqrt (x) , use the chain rule to find dy /dx when x =4

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{dy}{du}\frac{du}{dx}\] \[y = \frac{u-3}{u+3}\] so you need the quotient rule for this one, get \[\frac{dy}{du}=\frac{6}{(u+3)^2}\]

OpenStudy (anonymous):

\[u=\frac{8}{\sqrt{x}}\] \[\frac{du}{dx}=\frac{-4}{\sqrt{x^3}}\]

OpenStudy (anonymous):

multiply them together to get your answer, get \[\frac{dy}{dx}=\frac{6}{(u+3)^2}\times \frac{-4}{\sqrt{x^3}}\]

OpenStudy (anonymous):

and finally replace y by \[\frac{8}{\sqrt{x}}\]

OpenStudy (anonymous):

*replace u, not y

OpenStudy (anonymous):

srry where does the -4 / sqrt (x^3) from?

OpenStudy (anonymous):

oh and it said when x = 4!!

OpenStudy (anonymous):

so replace x by 4, or do it before multiplying

OpenStudy (anonymous):

\[\frac{8}{\sqrt{x}}=8x^{-\frac{1}{2}}\] now power rule

OpenStudy (anonymous):

get \[-4x^{-\frac{3}{2}}=\frac{-4}{\sqrt{x^3}}\]

OpenStudy (anonymous):

ohh ok i get it thanks!

OpenStudy (anonymous):

we can do this easier if we let x = 4 earlier on

OpenStudy (anonymous):

x = 4, \[\frac{8}{\sqrt{4}}=\frac{8}{2}=4\] so we can replace u by 4 in the derivative and get \[\frac{dy}{du}=\frac{6}{(4+3)^2}\]

OpenStudy (anonymous):

but it is not necessary

OpenStudy (anonymous):

oh ok thanks for your help :)